OpenStudy (anonymous):

Simplify √72/54?? PLEASE HELP I'LL AWARD AND FAN!!!

OpenStudy (studygurl14):

Is it like this?\(\Large\sqrt{\frac{72}{54}}\)

OpenStudy (studygurl14):

Or like this? \(\Large\frac{\sqrt{72}}{54}\)

OpenStudy (anonymous):

It's the 1st one :)

OpenStudy (studygurl14):

Okay, so that simplifies to \(\Large\frac{\sqrt{72}}{\sqrt{54}}\) To simplify the radicals, you need to factor the numbers underneath.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so divide by 2?

OpenStudy (studygurl14):

Yes.

OpenStudy (studygurl14):

What is 72 divided by 2?

OpenStudy (anonymous):

23

OpenStudy (studygurl14):

no.

OpenStudy (anonymous):

i meant 36 oops sorry

OpenStudy (anonymous):

and 54 divided by 2 is 27

OpenStudy (studygurl14):

yes, Okay, so you get \(\Large\frac{\sqrt{36\times2}}{\sqrt{27\times2}}\rightarrow\frac{\sqrt{36}\sqrt{2}}{\sqrt{27}\sqrt{2}}\rightarrow?\) What does \(\sqrt{36}\) equal? And you need to factor the 27 now

OpenStudy (anonymous):

27 is factored to 9, which can be factored to 3

OpenStudy (anonymous):

is that it?

OpenStudy (studygurl14):

Sort of. You get \(\Large \frac{\sqrt{36}\sqrt{2}}{\sqrt{9}\sqrt{3}\sqrt{2}}\) You need to simplify it now.

OpenStudy (anonymous):

square root 8 over 7??

OpenStudy (anonymous):

sorry this is all very confusing for me

OpenStudy (studygurl14):

Okay, let's take it one step at a time. What is \(\Large \sqrt{36}\) equal to?

OpenStudy (anonymous):

square root 3?

OpenStudy (studygurl14):

No. The square root of 36 is 6, remember? What is the square root of 9?

OpenStudy (anonymous):

ooh, the square root of 9 is 3

OpenStudy (anonymous):

so whats the next step?

OpenStudy (studygurl14):

right. So that simplifies to \(\Large\frac{6\sqrt{2}}{3\sqrt{3}\sqrt{2}}\) Now, you need to combine the \(\large\sqrt{3}\) and \(\large\sqrt{2}\), and simplify the 6/3

OpenStudy (anonymous):

would that be radical 5 and then 2/3

OpenStudy (anonymous):

is that okay??

OpenStudy (studygurl14):

Uh, no. It would be \(\Large \sqrt{3}\sqrt{2}=\sqrt{3\times 2}=\sqrt{6}\) And 6 divided by 3 is not 2/3

OpenStudy (anonymous):

oh, 2 then

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (anonymous):

are you still there?

OpenStudy (studygurl14):

Yes, so the final answer is \(\Large\frac{2\sqrt{2}}{\sqrt{6}}\) Unless your teacher doesn't want a radical in the denominator, in which case, we have more work to do